OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2 +bx+c=0. Solution: 1, only solution

OpenStudy (saifoo.khan):

now as the solution is one: (x=1) (x=1) (x-1)(x-1) = 0 multiply: x^2 -2x + 1 = 0

OpenStudy (anonymous):

saifoo khan speeding up the levels!

OpenStudy (saifoo.khan):

hehe, 77!!